# ISP-AD: A Large-Scale Real-World Dataset for Advancing Industrial Anomaly Detection with Synthetic and Real Defects

### v2

@@ 4.2 Training Settings
[…]
@@ 4.3.5 Implementation Details
To standardize the evaluation workflow, the Anomalib API [1] was utilized for the following unsupervised methods [64, 6]. Therefore, a training and evaluation pipeline, was implemented. Default method settings, as documented in the model configuration files, were applied. In case of Efficient-AD, the medium (M) patch descriptor network was chosen.
Additionally, the official implementations of DDAD [48] and GLASS [14], along with their configuration files, were selected. The target image conditioning parameters $w$ and $w\_DA$ were set to 3, together with 4 domain adaptation epochs, as specified in DDAD. In GLASS, the manifold hypothesis was selected as the GAS strategy, while foreground masks were omitted in LAS.
In general, the training input shape for image patches was set to 256 x 256 px. Despite center cropping (224 x 224 px) in DDAD for LSM-1, and center cropping (224 x 224 px) in ASM, no additional preprocessing was applied to the raw image patches. Due to Efficient-AD’s autoencoder architecture, the input shape was kept constant at 256 x 256 px. To enable feature extraction by means of the multi-scale vision transformer backbone in U-Flow, patches were resized to 448 x 448 px. Configuration files for all methods are publicly accessible in the dataset repository (Future Applications).
@@ 4.3.6 Training Settings
[…]
@@ 5 Experiments and Results
Table 2: Defect detection performance of the investigated mixed supervised training approach [26] evaluated on LSM-1. Starting from purely synthetic training, real defective patches were progressively injected with a probability of 1/32 at varying fractions (1/16, 1/8, 1/4, 1/2, 1) of collected defects. The best-performing fraction, based on image-level MCC, is highlighted in bold.
@@
Table 3: Defect detection performance of the investigated mixed supervised training approach [26] evaluated on ASM. Starting from purely synthetic training, real defective patches were progressively injected with a probability of 1/32 at varying fractions (1/16, 1/8, 1/4, 1/2, 1) of collected defects. The best-performing fraction, based on image-level MCC, is highlighted in bold.
@@
5.1 Defect Detection Performance Metrics
5.2 Mixed Supervised Training
To evaluate the scalability of the introduced mixed supervised training, different fractions of real defects accumulated during production were injected into the balanced stream of fault-free and synthetically generated defects.
Following random shuffling of all available defects, fractions of 1/2, 1/4, 1/8, and 1/16 were extracted in an ordered manner. This approach ensures that defects included in smaller fractions are also represented in larger ones. These fractions consist of both defect groups, points (Fig. 1 ) and area (Fig. 2), representing features that are either punctual or distributed across large sample areas. Additionally, to examine adaptability to unseen features, such as those missed during synthetization, training was performed exclusively on each defect group. In all experiments, starting from purely synthetic training, real defects were injected with a probability of 1/32.
To ensure comparability between experiments, the same training settings, including learning rates, were applied within each optical modality. Tables 2 and 3 visualize the image-level defect detection performance in LSM-1 and ASM.
5.2.1 Results
@@
Training with purely synthetic defects in LSM-1 resulted in an MCC of 0.81 with an FPR of 0.7 %. As described in the previous publication [35], depending on the underlying preprocessing strategy, FPRs less than 0.59 resp 0.33 % were necessary to avoid unwanted false alarms, leading to unnecessary sample rejections.
@@
Gradually adding real defective patches showed an increase in performance, achieving an MCC of 0.96 when utilizing all 172 available defects. This reduced the number of FN predictions from 22 to 4, representing a reduction of 81 %. By leveraging only 10 defective patches from both defect groups (mixed), an MCC of 0.91 was achieved. For comparison, supervised training with oversampling using all available (172) real defects, as described in [35], reached a comparable performance.
@@
Training with separated fractions of point and area defects also showed an increase in detection performance, with area defects having more impact. Adding this defect group reduced group-related FNs from 6 in purely synthetic training to 1 in both mixed training and area-specific fractions. In mixed fractions greater than 1/8, no area defect was overlooked.
@@
Training on the LSM-2 modality achieved near-perfect detection performance, overlooking only one defective patch. The injection of real defective patches resulted in faster convergence, requiring only 10 epochs compared to approximately 20 epochs for purely synthetic training.
@@
The ASM modality can be defined as the most demanding dataset due to its specific illumination characteristics (see Fig. 2) and imperfect segmentation masks. Training on purely synthetic defects resulted in an MCC of 0.70, which could be significantly improved to greater than 0.95 by utilizing fractions upwards of 1/8. This improvement reduced FNs by up to 93 % through the injection of real defects. The FPRs were comparable to those observed in LSM-1, indicating the method’s robustness to permitted sample variations and challenging inspection conditions.
@@
Leveraging the area defect group or mixed fractions had a significantly greater impact on improving detection performance compared to relying solely on punctual defects. In purely synthetic training, 23 out of 28 FNs were area defects. Training with injected area defects led to only a few overlooked (up to 3). Utilizing all 146 available defects resulted in just 2 FNs, one of each defect group.
@@
As described earlier, the capability to synthesize area defects is limited. In the case of ASM, defect synthetization for the squeegee stroke defect class (Fig. 2) was not possible and can be characterized by having unseen feature distributions. In contrast, LSM-1 contains grid defects (Fig. 2) that exhibit similar punctual features to the synthesized ones. As already observed in [35], LSM-1 demonstrate greater generalization capability using synthetic defects, though ASM benefits more from the injection of area defects. For the ASM modality, the injection of 9 area defects led to an increase in the MCC by 0.21, reaching 0.91. The injection of similar synthesized feature distributions, as in the case of point defects, proves beneficial for the overall defect detection performance, including previously overlooked punctual defects.
@@
Furthermore, training on several tens of thousands of fault-free patches introduced permitted normal feature variability, resulting in industry-applicable FPRs of less than 0.59 resp 0.33 %.
5.2.2 Limitations
@@
The above results are based on the initial random selection of defective samples and are therefore expected to vary with different seeds during dataset generation. Extensive hyperparameter tuning and statistical analysis are left for future research. In addition, the distribution of available area and points defect groups is imbalanced, due to reasons of availablility (Table 7).
@@ 5.3 Unsupervised Methods
To estimate the defect detection performance of unsupervised methods in comparison to the supervised approach described above, evaluations were conducted using the same imbalanced test datasets. In addition to image-level labels, ground truth masks were utilized to assess the segmentation performance of each method. Performance metrics of various SOTA unsupervised approaches, applied to modalities: LSM-1, LSM-2 and ASM, are summarized in Tables 4 to 6.
@@ 5.3.1 Results
Image-level defect detection performance in LSM-1 across different methods shows substantial variation with MCCs ranging from 0.60 at an FPR of 2.4 % to 0.89 at an FPR of 1.0 %. At the pixel level, minimum PRO-scores ranged from 4.1 to 28.3 %. In general, the low PRO-scores highlight the challenging task of detecting small punctual anomalies such as grid defects in Fig. 2. Common FPs occur at masked border regions or within areas of permitted variance in patterned structures (Fig. 4).
@@
The comparison of high pixel-level AUROCs ($>$ 99 %) compared to the measured PRO-scores illustrates their misleading interpretability in imbalanced scenarios.
Efficient-AD demonstrated a 7-fold reduction of FPR, compared to the reconstruction based approach (DDAD) based on denoising diffusion. Thus, Efficient-AD demonstrated an industry applicable FPR of 0.3 %, with qualitatively precise anomaly segmentations of small pronounced defects. However, despite the patch descriptor’s well-defined receptive field ensuring robustness, small defects remain boundary cases and are prone to misclassification. This is likely due to the bias introduced by ImageNet pretraining and the shallow patch descriptor architecture, which may limit its ability to extract fine-grained, in-domain features for subtle anomalies. In general, small and weakly contrasted point defects remain challenging even in GLASS, resulting in an recall of 92.6 %.
Compared to modality LSM-1, defect classes such as pinholes and pattern misalignments, which stand out prominently from the background pattern, exhibit more pronounced defect features. Consequently, both image- and pixel-level performance of various unsupervised approaches show improvements, achieving MCCs of up to 0.96 and PRO scores above 99 %.
@@
ASM consists of numerous patches containing imperfect masked image borders, due to varying imaging conditions and subsequent influence on segmentation performance of applied deep learning-based models. Additionally, this modality is characterized by large normal feature variations caused by the underlying imaging procedure, which relies on direct reflection.

### v3

@@ 4.2 Training Settings
The network training was performed using the stochastic gradient descent optimizer with the following parameters: a learning rate ranging from $1\text{\times}{10}^{-5}\text{\,}\mathrm{t}$o $5\text{\times}{10}^{-4}\text{\,}\mathrm{,}$ weight decay of $1\text{\times}{10}^{-2}\text{\,}\mathrm{,}$ and momentum of 0.9. Cosine annealing with warm restarts ($To=7813$, T_mult=2T\_mult=2) was applied as the learning rate scheduler, following the approach outlined in [50]. Based on a backbone ablation experiment comparing ResNet [29], EfficientNet [72], and ConvNeXt [47], ResNet-18 pre-trained on ImageNet [35] was selected as the backbone, as it provided the best accuracy–efficiency trade-off for the underlying use case (see Section 5.2.1). The final fully connected layer, originally consisting of 1000 output neurons (as per ImageNet pretraining), was replaced by a layer with two output neurons representing the binary class labels ‘good’ and ‘defective’. To mitigate potential domain bias from pretrained weights, all layers of the selected architecture were fine-tuned using the initial learning rate and scheduling scheme described above. As additional regularization, dropout with a rate of $10\%$ was applied to the final fully connected layer of the classifier head.
As part of the preprocessing, patches from LSM-1 were brightness-adjusted by a factor of 1.5. Additionally, weak smoothing was applied across all modalities using a Gaussian kernel with a kernel size of 3 and a sigma of 1. In mixed training scenarios using both synthetic and real defects, the injection probability $p_{\text{inj}}$ for real defects was set to $1/32$. Thus, for a batch size of 128, the expected number of augmented real defects per batch was four. This setting aims to prevent overrepresentation of real defects within the balanced synthetic data stream.
Mixed precision training was carried out until the validation loss showed no improvement for 20 consecutive epochs. Model selection was based on the lowest validation loss. As previously described, validation was performed on the available test set. For the binary classification task, thresholds were determined according to the optimal F1-scores obtained.
@@ 4.3.5 Implementation Details
To standardize the evaluation workflow, the Anomalib API [1] was utilized for the following unsupervised methods [71, 7]. Therefore, a training and evaluation pipeline, was implemented. Default method settings, as documented in the model configuration files, were applied. In case of Efficient-AD [7], the medium (M) patch descriptor network was chosen.
Additionally, the official implementations of DDAD [53] and GLASS [15], along with their configuration files, were selected. The target image conditioning parameters $w$ and w_DAw\_DA were set to 3, together with 4 domain adaptation epochs, as specified in DDAD. In GLASS, the manifold hypothesis was selected as the GAS strategy, while foreground masks were omitted in LAS.
In general, the training input shape for image patches was set to 256 x 256 px. Despite center cropping (224 x 224 px) in DDAD for LSM-1, and center cropping (224 x 224 px) in ASM, no additional preprocessing was applied to the raw image patches. Center cropping within ASM excludes any masked borders produced by manual patch extraction (Section 3.2.5). Due to Efficient-AD’s autoencoder architecture, the input shape was kept constant at 256 x 256 px, however during evaluation, border regions were omitted, as shown in Fig. 4. To enable feature extraction by means of the multi-scale vision transformer backbone in U-Flow, patches were resized to 448 x 448 px. Configuration files for all methods are publicly accessible in the dataset repository (Future Applications).
@@ 4.3.6 Training Settings
Anomalib’s engine was trained until the pixel-level AUROC showed no improvement for 30 consecutive epochs, with a limit of 300 epochs. The denoising diffusion approach was trained for 3000 steps with learning rates ranging from $5\text{\times}{10}^{-5}\text{\,}\mathrm{t}$o $1\text{\times}{10}^{-4}\text{\,}$ During testing, the starting point for the denoising trajectory was set to 250, with a step size of 25. Due to memory limitations, the batch size for both training and testing was reduced to 8. GLASS was trained up to 160 meta-epochs, with learning rates ranging from $1.25\text{\times}{10}^{-5}\text{\,}\mathrm{t}$o $5\text{\times}{10}^{-5}\text{\,}\mathrm{,}$ also using a batch size of 8. Thresholds for both image- and pixel-level metrics were determined by computing the optimal F1-score at the image-level. Training was performed on reduced splits of up to 500 patches.
@@ 5 Experiments and Results
Table 2: Ablation study comparing oversampling [38] and mixed training ($p_{\text{inj}}=1/32$). [28] across different backbones (ResNet [29], EfficientNet [72], ConvNeXt [47]), different fractions of real defects, and a synthetic-only baseline on LSM-1. Metrics: MCC, inference time per patch, and throughput (mean $\pm$ std, patches/s). The best MCC in the full-data regime (172 real defects) is highlighted in bold, while the best MCC in the few-shot regime (10 real defects) is marked with a dagger ($\dagger$).
Model Strategy # Real Defects MCC Inf. Time (ms) Throughput (patches/s)
ResNet-18 Oversampling 10 $0.70$ $0.16$ $6309\pm 5.$
172 $0.91$
Mixed 0 $0.81$
10 $0.91$
172 0.96
ResNet-50 Oversampling 10 $0.78$ $0.45$ $2216\pm 1.$
172 $0.93$
Mixed 0 $0.74$
10 $0.91$
172 0.97
EfficientNet-B0 Oversampling 10 $0.73$ $0.22$ $4608\pm 5.$
172 $0.94$
Mixed 0 $0.86$
10 $0.91$
172 0.96
EfficientNet-B4 Oversampling 10 $0.74$ $0.55$ $1815\pm 1.$
172 0.98
Mixed 0 $0.87$
10 $0.94$
172 $0.97$
ConvNeXt-Tiny Oversampling 10 $0.80$ $0.52$ $1912\pm 9.$
172 0.98
Mixed 0 $0.92$
10 $0.96$
172 0.98
5.1 Defect Detection Performance Metrics
5.2 Mixed Supervised Training
To evaluate the scalability of the introduced mixed supervised training, different fractions of real defects accumulated during production were injected into the balanced stream of fault-free and synthetically generated defects. Additionally, to examine adaptability to unseen features, such as those missed during synthetization, training was performed exclusively on each defect group (points or area).
Following random shuffling of all available defects, fractions of 1/2, 1/4, 1/8, and 1/16 were extracted in an ordered manner. This approach ensures that defects included in smaller fractions are also represented in larger ones. These fractions consist of both defect groups, points (Fig. 1 ) and area (Fig. 2), representing features that are either punctual or distributed across large sample areas (mixed group). In all experiments, starting from purely synthetic training, real defects were injected with a probability $p_{\text{inj}}$ of 1/32. To ensure comparability between experiments, the same training settings, including learning rates and seeds, were applied within each optical modality.
Prior to investigating the scalability and generalizability of mixed supervised training, ablation experiments were conducted on LSM-1 to identify an industry-applicable backbone that satisfies process-cycle constraints defined in [38]. The results, summarized in Table 2, compare the oversampling approach proposed in [38] with mixed training [28] across different backbones, defect fractions, and synthetic-only baselines, thereby highlighting accuracy–efficiency trade-offs relevant for industrial inspection. Following the backbone selection, image-level defect detection performance in LSM-1 and ASM across various defect fractions is presented in Tables 3 and 4.
5.2.1 Ablation: Backbone Selection
As defined in [38], the overall inspection time, utilizing all three measurement chambers, was restricted to 15–30 s. To avoid bottlenecks induced by the inference stage, efficient backbones from standard ResNet architectures [29] as well as more recent EfficientNet and ConvNeXt architectures [72, 47] were evaluated.
The inference time measurements reported in Table 2 were determined as the arithmetic mean of five cycles with 100 repetitions, including preprocessing of the respective methods (GeForce RTX 4080 setup as described in Section 4.3.7). Each cycle was preceded by a 10-iteration warm-up. To account for asynchronous CUDA processing, inference times were measured using PyTorch’s synchronized CUDA events at a batch size of 128, with autocast applied for mixed-precision evaluation.
As shown in Table 2, all architectures achieved fast inference times below 1 ms per patch, although the shallow ResNet-18 outperformed the other backbones by a factor of 1.36 compared to EfficientNet-B0 and up to approximately 3.5 compared to EfficientNet-B4 and ConvNeXt-Tiny.
ConvNeXt-Tiny, a modernized CNN architecture that incorporates depthwise convolutions with larger $7\times 7$ kernels together with other design elements inspired by Vision Transformers [47], achieved the highest MCC values across most training strategies. With all available real defects, pure oversampling achieved a maximum MCC of 0.98. EfficientNet-B4 trained with all available defects also reached an MCC of 0.98. Under these full-data conditions, the mixed training strategy achieved a comparable level of performance.
For all other architectures, mixed training with all available defects yielded higher MCC values, particularly in standard backbones such as ResNet [29], achieving 0.96 (ResNet-18) and 0.97 (ResNet-50), compared to 0.91 and 0.93 with pure oversampling.
In the few-shot regime (10 real defects), pure oversampling was outperformed by the mixed training strategy by large margins, with MCC improvements of 0.16 (ConvNeXt-Tiny) to 0.21 (ResNet-18). Across all backbones, starting from the purely synthetic baseline, the injection of real defects into the balanced synthetic stream progressively improved defect detection. Even a few injected real defects proved beneficial, achieving MCC values greater than 0.91 across all backbones. Moreover, the ConvNeXt-Tiny architecture reached a high MCC of 0.96 with just 10 injected real defects, highlighting the superiority of the mixed strategy in few-shot regimes.
Furthermore, the synthetic-only baseline, which did not rely on any real defects for defect generation (model-free approach), outperformed few-shot (10 real defects) oversampling in 4 out of 5 backbone experiments.
Based on the given industrial requirements and to avoid any inference bottleneck, the best accuracy–efficiency trade-off was achieved with the ResNet-18 architecture, reaching performance close to that of more modern architectures when using the mixed training strategy on all available data. Nonetheless, depending on future industrial requirements, ConvNeXt architectures remain a viable choice, delivering industry-applicable detection performance even in the few-shot regime.
@@ 5.2.1 Ablation: Backbone Selection
Table 3: Defect detection performance of the investigated mixed supervised training strategy [28] evaluated on LSM-1. Starting from purely synthetic training, real defective patches were progressively injected with an injection probability $p_{\text{inj}}$ of 1/32 at varying fractions (1/16, 1/8, 1/4, 1/2, 1) of collected defects. The best-performing fraction, based on image-level MCC, is highlighted in bold.
@@
Table 4: Defect detection performance of the investigated mixed supervised training strategy [28] evaluated on ASM. Starting from purely synthetic training, real defective patches were progressively injected with an injection probability $p_{\text{inj}}$ of 1/32 at varying fractions (1/16, 1/8, 1/4, 1/2, 1) of collected defects. The best-performing fraction, based on image-level MCC, is highlighted in bold.
@@
5.2.2 Scalability and Synthetic Defect Complementation
Based on the ablation, ResNet-18 was selected for the subsequent experiments, as it represented the best accuracy-efficiency trade-off. The scalability of mixed training was evaluated on LSM-1, LSM-2, and ASM. Experiments examined whether small numbers of real defects with previously unseen characteristics—such as squeegee strokes in ASM or grid defects in LSM-1, which could not be synthesized—could complement synthetic features and improve generalization.
@@ 5.2.2 Scalability and Synthetic Defect Complementation
Training with purely synthetic defects in LSM-1 resulted in an MCC of 0.81 with an FPR of 0.7 %. As described in the previous publication [38], depending on the underlying preprocessing strategy, FPRs less than 0.59 respectively 0.33 % were necessary to avoid unwanted false alarms, leading to unnecessary sample rejections.
@@
Gradually adding real defective patches increased performance, achieving an MCC of 0.96 when utilizing all 172 available defects. This reduced the number of FN predictions from 22 to 4, representing an 81% reduction. Using only 10 defective patches from both defect groups (mixed) resulted in an MCC of 0.91. For comparison, supervised training with oversampling on all 172 available real defects achieved the same MCC of 0.91 (Table 2).
@@
Training with separated fractions of point and area defects also showed improved detection performance, with area defects having more impact. Adding this defect group reduced group-related FNs from 6 in purely synthetic training to 1 in both mixed training and area-specific fractions. In mixed fractions greater than 1/8, no area defect was overlooked.
@@
Training on the LSM-2 modality achieved near-perfect detection performance, overlooking only a single defective patch. The injection of real defective patches accelerated convergence, requiring only 10 epochs compared to approximately 20 epochs for purely synthetic training. This behavior may be attributed to the clearly pronounced defect characteristics of pinholes and pattern misalignment, as illustrated in Fig. 1 and Fig. 2.
@@
The ASM modality can be defined as the most demanding dataset due to its specific illumination characteristics (see Fig. 2) and imperfect segmentation masks. Training on purely synthetic defects resulted in an MCC of 0.70, which could be significantly improved to greater than 0.95 by utilizing fractions upwards of 1/8. This improvement reduced FNs by up to 93 % through the injection of real defects. The FPRs were comparable to those observed in LSM-1, indicating robustness to permitted sample variations and challenging inspection conditions.
@@
Leveraging the area defect group or mixed fractions had a significantly greater impact on detection performance compared to relying solely on punctual defects. In purely synthetic training, 23 out of 28 FNs were area defects. Training with injected area defects reduced this number to only a few (up to 3). Utilizing all 146 available defects resulted in just 2 FNs, one of each defect group.
Additional experiments using solely real defects within the supervised oversampling approach [38] achieved MCCs greater than 0.98 for fractions upwards of 1/4. The more distinct defect characteristics compared to LSM-1 (e.g. scratches or dots), as well as the limited defect diversity within the test set (see Section 3.2.5), might be reasonable explanations for observed performance saturation.
As observed across other backbones in LSM-1 (Table 2), the performance of the proposed mixed training strategy remained superior in the low data regimes (fractions 1/16 and 1/8), especially when training was performed exclusively on area and point defect groups. In these settings, MCCs of 0.91 and 0.76 were achieved, respectively (see Table 4), compared to MCCs of 0.56 and 0.65 for pure oversampling. Furthermore, optimized F1 thresholds below 0.1 indicated majority-class bias in the supervised oversampling strategy in these few-shot regimes.
Complementation with Real Defects
@@
As described earlier, the capability to synthesize area defects is limited. In the case of ASM, defect synthetization for the squeegee stroke defect class (Fig. 2) was not possible and can be characterized by having unseen feature distributions. In contrast, LSM-1 contains grid defects (Fig. 2) that exhibit similar punctual features to the synthesized ones. Models trained with synthetic defects generalized better on LSM-1, whereas ASM benefited more strongly from the injection of area defects.
@@
For the ASM modality, the injection of 9 area defects led to an increase in the MCC by 0.21, reaching 0.91. The injection of similar synthesized feature distributions, as in the case of point defects, proved beneficial for the overall defect detection performance, including previously overlooked punctual defects. Furthermore, training on several tens of thousands of fault-free patches introduced permitted normal feature variability, resulting in industry-applicable FPRs of less than 0.59 respectively 0.33 %.
5.2.3 Key Findings Mixed Training Strategy
Above experiments indicate that mixed training achieves scalable performance across all investigated modalities, with synthetic and real defects complementing each other. (i) In LSM-1, model-free synthetization established a strong initial boundary, achieving an MCC of up to 0.92 in the synthetic-only baseline with ConvNeXt-Tiny. Its capability, however, appeared limited by the synthetization method and the underlying modality: in LSM-1, punctual defect characteristics such as those occurring in grid defects could be captured to a significant degree synthetically, whereas in ASM, unsynthesizable morphologies such as squeegee strokes required real injection. (ii) Injecting even a few real defects complemented the already learned synthetic features, reducing false negatives and refining the initial boundary. (iii) Mixed training consistently showed superior performance in few-shot regimes, achieving an MCC of at least 0.91 with only 10 real defect samples across various backbones when evaluated on LSM-1. It was beneficial in avoiding majority-class bias, converged faster, and maintained the low FPRs required in industrial inspection. The mixed training strategy proved architecture-agnostic, improving results across ResNet [29], EfficientNet [72], and ConvNeXt [47], with clear trade-offs between detection performance and inference speed.
Together, these findings suggest that synthetic defects can serve as a cold-start baseline, while injected real defects provide complementary information for defect characteristics that cannot be synthesized. Furthermore, these results demonstrate that synthetic data can substantially reduce the number of real samples needed to achieve a given level of detection performance.
Thus, model-free synthesis approaches, as implemented in the proposed algorithm [27], which does not require defective training data for elaborate model training, may represent a viable option for rapid deployment in industry, as demonstrated in the proposed industrial use case [38].
5.2.4 Limitations and Future Adaptations
@@
The above results are based on the initial random selection of defective samples and are therefore expected to vary with different seeds during dataset generation. Extensive hyperparameter tuning and statistical analysis are left for future research. In addition, the distribution of available "area" and "points" defect groups is imbalanced, due to limited availability (Table 7).
While this approach avoids the need for a trained generative model, hyperparameter tuning of the defect synthesis algorithm can be elaborate, and the diversity of synthesized defects remains limited.
Depending on data availability and deployment stage, more recent approaches such as generative models or vision language models [80] could be integrated as synthesis modules within this adaptable closed-set approach. A combination of model-free approaches (e.g. based on Perlin masks [55]) with model-based approaches (e.g. diffusion-based models [26]) could further enhance feature diversity. ISP-AD therefore provides a benchmark, with its additional fault-free data splits enabling the investigation of emerging synthetization methods (see Section Declarations).
@@ 5.3 Unsupervised Methods
To estimate the defect detection performance of unsupervised methods in comparison to the supervised approach described above, evaluations were conducted using the same imbalanced test datasets. In addition to image-level labels, ground truth masks were utilized to assess the segmentation performance of each method. Performance metrics of various SOTA unsupervised approaches, applied to modalities: LSM-1, LSM-2 and ASM, are summarized in Tables 5 to 7.
@@ 5.3.1 Results
Image-level defect detection performance in LSM-1 across different methods shows substantial variation with MCCs ranging from 0.60 at an FPR of 2.4 % to 0.89 at an FPR of 1.0 %. At the pixel level, minimum PRO-scores ranged from 4.1 to 28.3 %. In general, the low PRO-scores highlight the challenging task of detecting small punctual anomalies such as grid defects in Fig. 2. Common FPs occur at masked border regions or within areas of permitted variations in patterned structures (Fig. 4).
@@
The comparison of high pixel-level AUROCs (>> 99 %) compared to the measured PRO-scores illustrates their misleading interpretability in imbalanced scenarios.
Efficient-AD demonstrated a 7-fold reduction of FPR, compared to the reconstruction based approach (DDAD) based on denoising diffusion. Thus, Efficient-AD demonstrated an industry-applicable FPR of 0.3 %, with qualitatively precise anomaly segmentations of small pronounced defects. However, despite the patch descriptor’s well-defined receptive field ensuring robustness, small defects remain boundary cases and are prone to misclassification. This is likely due to the bias introduced by ImageNet pretraining and the shallow patch descriptor architecture, which may limit its ability to extract fine-grained, in-domain features for subtle anomalies. In general, small and weakly contrasted point defects remain challenging even in GLASS, resulting in an recall of 92.6 %.
Compared to modality LSM-1, defect classes such as pinholes and pattern misalignments in LSM-2 (Fig. 1 and Fig. 2), which stand out prominently from the background pattern, exhibit more pronounced defect features. Consequently, both image- and pixel-level performance of various unsupervised approaches show improvements, achieving MCCs of up to 0.96 and PRO scores above 99 %.
@@
ASM consists of numerous patches containing imperfect masked image borders, due to varying imaging conditions and the subsequent influence on segmentation performance of the applied deep learning-based model. These artifacts reflect practical deployment conditions and were intentionally retained in the dataset to simulate realistic segmentation imperfections. Additionally, this modality is characterized by large normal feature variations caused by the underlying imaging procedure, which relies on direct reflection.